OpenStudy (heart_offire):

Simplify. x^2 + 5x - 14 ----------- x^2 + 8x + 7 A) 7x -- x B) x - 2 ---- x + 1 C) x - 2 ---- x + 7 D) x + 7 ----- x + 1 @study312 i feel like its D

OpenStudy (study312):

yeah

OpenStudy (inkyvoyd):

yeah... no

OpenStudy (inkyvoyd):

factor the numerator and denominator... what do you get

OpenStudy (heart_offire):

which numerator and denominator

OpenStudy (inkyvoyd):

the original problem

jhonyy9 (jhonyy9):

so you can factorizing x^2 +5x-14 = (x+7)(x-2) - yes ? than now try the second equation

jhonyy9 (jhonyy9):

so x^2 +8x+7 = x*x +x+7x +7 = x(x+1) +7(x+1) = (x+1)(x+7) yes ? hope this will help you and you can ending your exercise

OpenStudy (heart_offire):

^thanks for the help :)